how many terms of the arithmetic series 10 + 8 + 6 +... will make the sum 24?



1
Expert's answer
2021-07-19T13:22:57-0400

This is an arithmetic sequence with "a_1=10, d=-2."

"S_n=\frac{(2a_1+(n-1)d)n}{2}=24."

"n^2-11n+24=0."

"n_1=\frac{11-\sqrt{121-96}}{2}=3."

"n_2=\frac{11+\sqrt{121-96}}{2}=8."

3 or 8 terms of this arithmetic series will make the sum 24.